# Fairness in Ranking under Disparate Uncertainty

### v3

@@ Abstract.
copyright: noneccs: Information systems Ranking
@@ 1. Introduction
Human decision-processes are increasingly augmented with algorithmic decision-support systems, which has created opportunities and challenges for addressing group-based disparities in decision outcomes (Dixon-Roman et al., 2013; Buolamwini and Gebru, 2018; Tatman, 2017; Wilson et al., 2019). In this paper, we focus on selection processes where humans evaluators use rankings to organize the order of review under resource constraints. We argue that disparities in uncertainty can be a major source of group-based discrimination in this setting.
@@
  • In addition to these theoretical worst-case guarantees, we present extensive experiments benchmarking the EOR algorithm with various existing ranking algorithms under different settings of disparate uncertainty. We show that Demographic Parity (Yang and Stoyanovich, 2016; Zehlike et al., 2017), normative procedures like Proportional Rooney-rule-like constraints (Celis et al., 2020), Exposure based fairness criteria (Singh and Joachims, 2018), and Thompson Sampling Policy (Singh et al., 2021) are not typically EOR-fair under disparate uncertainty. We find that these results hold on both a wide range of synthetic datasets, as well as on real-world US census data. Finally, we explore the use of our fairness criterion for auditing ranking systems, using a real-world dataset of Amazon shopping search queries. Our code can be accessed at https://github.com/RichRast/Fairness_in_Ranking_under_Disparate_Uncertainty.
@@ 2. Related Works
Fairness in Rankings and Selection Processes: While there exist several notions of fairness in rankings (Zehlike et al., 2021), predominantly, they are variations of two fairness mechanisms for rankings in existing literature – representation by size (Celis et al., 2017; Yang et al., 2019; Zehlike et al., 2017, 2022) and equitable allocation of exposure (Singh and Joachims, 2018, 2017; Biega et al., 2018; Konstantinov and Lampert, 2021; Narasimhan et al., 2020). We propose a new criterion different from either of the two and our central point is that under disparate uncertainty between groups, it is more fair to take equal amounts of relevance in expectation rather than equality by size or exposure. Proportional representation in the form of diversity constraints like demographic parity (Yang and Stoyanovich, 2016) or affirmative action such as the Rooney Rule (Celis et al., 2020) guarantee a minimum proportion by group size in selection processes. Exposure based formulations in rankings ensure that groups of candidates are allocated exposure in an equitable way such as in proportion of amortized relevance over the full ranking (Biega et al., 2018). In this work, we demonstrate that fairness of representation by size and exposure, are not sufficient under disparate uncertainty.
Fairness in Rankings under Uncertainty: Our work builds on (Singh et al., 2021), in which the authors establish that uncertainty in relevance probabilities is a primary cause of unfairness for rankings. They propose a Thompson sampling policy that randomizes relevances drawn from the predictive posterior distribution. Separately, (Emelianov et al., 2022) studies the role of affirmative action in the presence of differential variance between groups in rankings. Differential variance implies that there is more certainty about the true quality (scores) of candidates in a group with less variance in the estimated quality and vice versa for a group with higher variance. In contrast, we work with relevance probabilities instead of scores and focus on the certainty of relevance of a candidate, which is determined by how close their relevance probabilities are to 1 or 0. For example, a group is highly certain (if the probabilities are all close to 1.0) or highly uncertain (if the probabilities are all close to 0.5) but both have similar variance in probabilities. Fairness under uncertainty has also been studied with respect to calibration of probabilities (Pleiss et al., 2017; Kleinberg et al., 2017; Flores et al., 2016; Chouldechova, 2017). Classical literature in this area studies whether group-wise calibration is a necessary condition for fairness, or not (Loi and Heitz, 2022). Our work is orthogonal to the question of the necessity of calibration for fairness and we only require group-wise calibration as a sufficient condition for the EOR criterion we propose.
@@ 3.1. Illustrative Example
We argue that a more principled and fair way would be to select an equal fraction of relevant candidates from each group in expectation. Consider the last ranking in Figure 2, which approximately fulfills the EOR criterion we formally introduce later. In expectation, this ranking selects a more equal number of relevant candidates from both groups, making it similar to a fair lottery. In particular, it selects 1.8 out of 4 relevant candidates from group A and 1.2 out of 4 relevant candidates from group B. This EOR ranking, however, comes at an increased evaluation cost to the principal as it selects 3.0 expected relevant candidates from both the groups, compared to 3.2 with DP and 3.3 with PRP. As a result, the principal needs to review more candidates to select the same number of relevant candidates with EOR ranking. However, it is still far more effective than a lottery, which selects the candidates in a uniform random order.
@@
This example illustrates the intuition behind the EOR principle we formalize in the following, and we will show how to efficiently compute rankings that fulfill our EOR criterion.
@@ 4. Equality of Opportunity in Ranking
$nRel(g)=\sum_{i\in g}\mathbb{E}_{r_{i}\sim(\mathbb{P}(r_{i}|\mathcal{D}))}[r_{i}]=\sum_{i\in g}\mathbb{P}(r_{i}=1|\mathcal{D})$
@@
$nRel(g|\sigma_{k})=\sum_{i\in g\cap\sigma_{k}}\mathbb{E}[r_{i}]=\sum_{i\in g\cap\sigma_{k}}\mathbb{P}(r_{i}=1|\mathcal{D}).$
@@
(1) $nRel(g|\pi_{k})$ $=$ $\sum_{i\in g}\mathbb{E}_{r_{i}\sim(\mathbb{P}(r_{i}|\mathcal{D})),\sigma_{k}\sim\pi}[r_{i}\mathbb{I}_{i\in\sigma_{k}}]$
$=$ $\sum_{i\in g}\mathbb{P}(i\in\sigma_{k}^{\pi})\mathbb{P}(r_{i}=1|\mathcal{D}),$
@@ 4.1. Cost Burden to Candidate Groups and the Principal
(3) $c(g|\pi_{k})$ $=$ $\frac{\sum_{i\in g}\mathbb{E}_{r_{i}\sim(\mathbb{P}(r_{i}|\mathcal{D}))}[c(i|\pi_{k},r_{i}=1)]}{nRel(g)}$
@@
The last equality in (3) follows directly from Equation (1). We normalize the expected group cost with the total expected number of relevant candidates in the group so that the above approximates the fraction of relevant candidates from that group that miss out on the opportunity of being selected by the human reviewers.
@@
We again normalize this quantity to make it proportional to the total expected number of relevant candidates. Note that (4) is related to the conventional metric of Recall@k.
Similarly, related to the Precision@k metric, there is a cost to the Principal quantified via the expected number of relevant candidates out of $k$ candidates given by $\frac{\sum_{i}(1-\mathbb{P}(i\in\sigma_{k}^{\pi}))\mathbb{P}(r_{i}=1|\mathcal{D})}{k}$. In this work, we focus on the cost burden for all stakeholders and leave a detailed analysis of Precision@k and Recall@k tradeoff for future work. In the rest of the paper, we analyze and discuss the cost burden to candidate groups and Principal as defined in Equation (3) and (4) respectively.
@@ 4.2. Equality of Opportunity Ranking (EOR) Criterion
property that the expected number of relevant candidates for each group in the top $k$ never differs by more than $0.6$ for any value of $k$. In one way, this guarantee is even stronger than what is defined in Axiom 1, since it holds for the specific ranking $\sigma^{EOR}$ without the need for stochasticity in the ranking policy. This provides a non-amortized notion of fairness, which is particularly desirable for high-stakes ranking tasks that do not repeat, and we thus need to provide the strongest possible guarantees for the specific ranking $\sigma$ we present. However, a guarantee for an individual ranking makes the problem inherently discrete, which means that we require some tolerance (i.e., $0.6$ in the example above) in the fairness criterion depending on the choice of $k$. This leads to the following $\delta$-EOR Fairness criterion for an individual ranking $\sigma$.
@@
$\delta$-EOR fairness balances the selection of candidates from the two groups, accounting for predictive uncertainty in their estimation of relevances. If for instance, the ML model is less certain in its predictions for group B, but both groups have the same total expected relevance, the EOR criterion will rank candidates from group B higher to ensure fairness. Importantly, note how this produces more human relevance labels of candidates from groups with high uncertainty, which has the desirable side-effect of producing new training data that allows training of more equitable relevance models for future use.
Finally, note how the $\delta$-EOR criterion provides a means for ensuring procedural fairness and avoiding disparate treatment. Importantly, we leave the decision of which candidates to select to the human decision maker, and the EOR criterion does not require the designation of a disadvantaged group. Instead, the EOR criterion is symmetrical w.r.t. both groups and by definition treats both groups similarly, and its intervention in the ranking process is entirely driven by the predictive model $\mathbb{P}(r_{i}|\mathcal{D})$. Even though it uses group membership, EOR-fairness is thus fundamentally different from demographic parity (Dwork et al., 2012; Yang and Stoyanovich, 2016) and affirmative action rules like Rooney rule (Collins, 2007; Celis et al., 2020), $\frac{4}{5}$th rule (selection rate for a protected group must be at least 80% of the rate for the group with the highest rate) [^3] Uniform Guidelines on Employment Selection Procedures, 29 C.F.R.§1607.4(D) (2015) or $\gamma$-based notions of fairness (Emelianov et al., 2020) and threshold based formulations such as FAIR (Zehlike et al., 2017).
@@
One of our key contributions includes formalizing the connection between $\delta$-EOR Fair Ranking described in 4.1 and the cost of opportunity in rankings described in Section 4.1. Both $\delta$-EOR Fair Ranking and cost of opportunity in rankings are derived separately – the former from the axiom of fairness of a uniform lottery, the latter from the cost of errors that any realistic prediction model is bound to make. In the next section, we show that these two are elegantly related via theoretical results on cost optimality.
@@ 5. Computing EOR-Fair Rankings
Initialize: $j\leftarrow 0$; empty ranking $\sigma^{EOR}$
1:while $j<k$ do
2:     $l_{g}\leftarrow\sigma^{PRP,g}[1]\hskip 10.0pt\forall g\in\{A,B\}$
3:     $g^{*}\leftarrow\underset{g\in\{A,B\}}{\arg\min}\left|\delta(\sigma^{EOR}\cup\{l_{g}\})\right|,\text{where }\delta(.)$
[…]
5:     $l_{g^{*}}\leftarrow\sigma^{PRP,g^{*}}[1];\hskip 5.0pt\sigma^{PRP,g^{*}}\leftarrow\sigma^{PRP,g^{*}}\textbackslash\{l_{g^{*}}\}$
6:     $\sigma^{EOR}$leftarrow $\sigma^{EOR}\cup\{l_{g^{*}}\}$;    $j\leftarrow j+1$
7:end while
8:return $\sigma^{EOR}$
While Algorithm 1 is inspired by existing algorithms such as (Zehlike et al., 2017) in that both select the top element from the PRP ranking of each group, they are fundamentally different. Existing methods including (Zehlike et al., 2017) ensure a form of demographic parity which we have already shown to be fundamentally different than the EOR criterion we propose. Additionally, while (Zehlike et al., 2017) requires a threshold input and the designation of a disadvantaged group, the EOR Algorithm does not require this normative designation and guarantees EOR fairness without requiring any tolerance $\delta$ as an input. We show this both theoretically and in empirical evaluations and provide a detailed description of baseline algorithms in Section E.1.
@@
Minimize total cost as defined in Equation 4
$\text{Minimize}_{x\in\{0,1\}}1-\frac{P^{T}X}{nRel(A)+nRel(B)}$
subject to $\text{subject to }X^{T}\mathbb{1}=k\hskip 5.0pt\text{(select up to $k$ candidates)}$)
$-\delta(\sigma^{EOR}_{k})\leq\left(\frac{P\mathbb{I}_{A}}{nRel(A)}-\frac{P\mathbb{I}_{B}}{nRel(B)}\right)^{T}X\leq\delta(\sigma^{EOR}_{k})$
(The EOR criterion Equation 6 must be satisfied $\forall k$)
We relax this ILP to a Linear Program (LP) by turning any integer constraints $x\in\{0,1\}$ in the primal into $0\leq x\leq 1$ For the relaxed LP, we formulate its dual and construct a set of dual variables $\lambda$ corresponding to the solution from the EOR Algorithm. Using the dual value of the EOR solution and the relaxed LP solution, we obtain an upper bound of the duality gap. Since the upper bound on this duality gap is w.r.t. the relaxed LP solution, it is also an upper bound for the optimal ILP solution. We provide a complete proof of the theorem and associated lemmas in Section C.1. ∎
@@
$\phi\delta_{max}=\frac{1}{nRel(A)+nRel(B)}\left|\frac{p_{A}-p_{B}}{q_{A}-q_{B}}\right|\left(\frac{\sigma^{PRP,A}[1]}{nRel(A)}+\frac{\sigma^{PRP,B}[1]}{nRel(B)}\right).$ Further, we show that if the EOR algorithm selects all the elements from one group at some position $k$, then selecting the remaining elements from the other group satisfies the $\delta_{max}$ constraint. We provide a complete proof of Theorem 5.2 in Section C.2. ∎
We now consider the comparison of EOR with the Uniform ranking policy and the Probability Ranking Principle. In Proposition 5.3, we focus the analysis on positions $k$ with $\delta=0$ to avoid discretization effects.
@@
In Proposition 5.4, we show that any difference in the costs between the utility maximizing PRP policy and the EOR policy is related to $\delta(\sigma)$ of the two rankings.
Proposition 5.0 (Costs from EOR vs. PRP).
Any difference in costs between EOR ranking and PRP ranking directly relates to the EOR criterion $\delta(\sigma_{k})$ of the two rankings as follows
$c(\text{Principal}|\sigma_{k}^{EOR})-c(\text{Principal}|\sigma_{k}^{PRP})$ $=$
$\frac{nRel(A)\delta(\sigma_{k}^{PRP})}{nRel(A)+nRel(B)}+\frac{nRel(B)\delta(\sigma_{k}^{EOR})}{nRel(A)+nRel(B)}$displaystyle-- $\delta(\sigma_{k}^{PRP}\cap\sigma_{k}^{EOR})$
if $nRel(A|\sigma_{k}^{PRP})\geq nRel(A|\sigma_{k}^{EOR})$.
We provide proof of proposition Proposition 5.3 and 5.4 in Section C.3 and Section C.4 respectively.
In summary, we have shown that Algorithm 1 is an efficient algorithm that computes rankings close to the optimal solution, making it a promising candidate for practical use.

### v4

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Equity and Access in Algorithms, Mechanisms, and Optimization; October 29–31, 2024; San Luis Potosi, Mexicobooktitle: Equity and Access in Algorithms, Mechanisms, and Optimization (EAAMO ’24), October 29–31, 2024, San Luis Potosi, Mexicodoi: 10.1145/3689904.3694703isbn: 979-8-4007-1222-7/24/10ccs: Information systems Rankings; Top-k retrieval; Recommender systems; Decision support systems
@@ 1. Introduction
Human decision-processes are increasingly augmented with algorithmic decision-support systems, which has created opportunities and challenges for addressing group-based disparities in decision outcomes. In this paper, we focus on selection processes where humans evaluators use rankings to organize the order of review under resource constraints. We argue that disparities in uncertainty can be a major source of group-based discrimination in this setting.
@@
  • In addition to these theoretical worst-case guarantees, we present extensive experiments benchmarking the EOR algorithm with various existing ranking algorithms under different settings of disparate uncertainty. We show that Demographic Parity (Yang and Stoyanovich, 2016; Zehlike et al., 2017), normative procedures like Proportional Rooney-rule-like constraints (Celis et al., 2020), Exposure based fairness criteria (Singh and Joachims, 2018), and Thompson Sampling Policy (Singh et al., 2021) are not typically EOR-fair under disparate uncertainty. We find that these results hold on both a wide range of synthetic datasets, as well as on real-world US census data. Finally, we explore the use of our fairness criterion for auditing ranking systems, using a real-world dataset of Amazon shopping search queries. Our code can be accessed at https://github.com/RichRast/DisparateUncertainty.
@@ 2. Related Works
Fairness in Rankings and Selection Processes: While there exist several notions of fairness in rankings (Zehlike et al., 2021), predominantly, they are variations of two fairness mechanisms in existing literature – representation by size (Celis et al., 2017; Yang et al., 2019; Zehlike et al., 2017, 2022) and equitable allocation of exposure (Singh and Joachims, 2018, 2017; Biega et al., 2018; Konstantinov and Lampert, 2021; Narasimhan et al., 2020). We propose a new criterion different from either of the two and our central point is that under disparate uncertainty between groups, it is more fair to take an equal proportion of relevance in expectation rather than equality by size or exposure. Proportional representation in the form of diversity constraints like demographic parity (Yang and Stoyanovich, 2016) or affirmative action such as the Rooney Rule (Celis et al., 2020) guarantee a minimum proportion by group size in selection processes. Exposure based formulations in rankings ensure that groups of candidates are allocated exposure in an equitable way such as in proportion of amortized relevance over the full ranking (Biega et al., 2018). In this work, we demonstrate that fairness of representation by size and exposure, are not sufficient under disparate uncertainty.
Fairness in Rankings under Uncertainty: Our work builds on (Singh et al., 2021), in which the authors establish that uncertainty in relevance probabilities is a primary cause of unfairness for rankings. They propose a Thompson sampling policy that randomizes relevances drawn from the predictive posterior distribution. Separately, (Emelianov et al., 2022) studies the role of affirmative action in the presence of differential variance between groups in rankings. Differential variance implies that there is more certainty about the true quality (scores) of candidates in a group with less variance in the estimated quality and vice versa for a group with higher variance. In contrast, we work with relevance probabilities instead of scores and focus on the certainty of relevance of a candidate, which is determined by how close the predicted relevance probabilities are to 1 or 0. For instance, a group is highly certain (if the probabilities are all close to 1.0) or highly uncertain (if the probabilities are all close to 0.5) while both groups could have similar variance in probabilities. Fairness under uncertainty has also been studied with respect to calibration of probabilities (Pleiss et al., 2017; Kleinberg et al., 2017; Flores et al., 2016; Chouldechova, 2017). Classical literature in this area studies whether group-wise calibration is a necessary condition for fairness, or not (Loi and Heitz, 2022). Our work is orthogonal to the question of the necessity of calibration for fairness and we only require group-wise calibration as a sufficient condition for the EOR criterion we propose.
@@ 3.1. Illustrative Example
We argue that a more principled and fair way would be to select an equal fraction of relevant candidates from each group in expectation. Consider the last ranking in Figure 2, which approximately fulfills the EOR fairness we formally introduce later. In expectation, this ranking selects a more equal number of relevant candidates from both groups, making it similar to a fair lottery. In particular, it selects 1.8 out of 4 relevant candidates from group A and 1.2 out of 4 relevant candidates from group B. This EOR ranking, however, comes at an increased evaluation cost to the principal as it selects 3.0 expected relevant candidates from both the groups, compared to 3.2 with DP and 3.3 with PRP. As a result, the principal needs to review more candidates to select the same number of relevant candidates with EOR ranking. However, it is still far more effective than a lottery, which selects the candidates in a uniform random order.
@@
This example illustrates the intuition behind the EOR principle we formalize in the following, and we will show how to efficiently compute rankings that fulfill EOR fairness.
@@ 4. Equality of Opportunity in Ranking
$nRel(g)=\sum_{i\in g}\mathbb{E}_{r_{i}\sim\mathbb{P}(r_{i}|\mathcal{D})}[r_{i}]=\sum_{i\in g}\mathbb{P}(r_{i}=1|\mathcal{D})$
@@
$nRel(g|\sigma_{k})=\sum_{i\in g\cap\sigma_{k}}\mathbb{E}[r_{i}]=\sum_{i\in g\cap\sigma_{k}}\mathbb{P}(r_{i}=1|\mathcal{D})$
@@
(1) $nRel(g|\pi_{k})$ $=$ $\sum_{i\in g}\mathbb{E}_{r_{i}\sim\mathbb{P}(r_{i}|\mathcal{D}),\sigma_{k}\sim\pi}[r_{i}\mathbb{I}_{i\in\sigma_{k}}]$
$=$ $\sum_{i\in g}\mathbb{P}(i\in\sigma_{k}^{\pi})\mathbb{P}(r_{i}=1|\mathcal{D})$
@@ 4.1. Cost Burden to Candidate Groups and the Principal
(3) $c(g|\pi_{k})$ $=$ $\frac{\sum_{i\in g}\mathbb{E}_{r_{i}\sim\mathbb{P}(r_{i}|\mathcal{D})}[c(i|\pi_{k},r_{i}=1)]}{nRel(g)}$
@@
The last equality in (3) follows directly from Eq. (1). We normalize the expected group cost with the total expected number of relevant candidates in the group so that the above approximates the fraction of relevant candidates from that group that miss out on the opportunity of being selected by the human reviewers.
@@
We again normalize this quantity to make it proportional to the total expected number of relevant candidates. Note that Eq. (4) is related to the conventional metric of Recall@k.
@@ 4.2. Equality of Opportunity Ranking (EOR) Criterion
$\delta$-EOR fairness balances the selection of candidates from the two groups, accounting for predictive uncertainty in their estimation of relevances. If for instance, the ML model is less certain in its predictions for group B, but both groups have the same total expected relevance, the $\delta$-EOR criterion will rank candidates from group B higher to ensure fairness. Importantly, note how this produces more human relevance labels of candidates from groups with high uncertainty, which has the desirable side-effect of producing new training data that allows training of more equitable relevance models for future use.
Finally, note how the $\delta$-EOR fair ranking provides a means for ensuring procedural fairness and avoiding disparate treatment. Importantly, we leave the decision of which candidates to select to the human decision maker, and EOR fairness does not require the designation of a disadvantaged group. Instead, the EOR fair condition in Eq. (6) is symmetrical w.r.t. both groups and by definition treats both groups similarly, and its intervention in the ranking process is entirely driven by the predictive model $\mathbb{P}(r_{i}|\mathcal{D})$. Even though it uses group membership, EOR-fairness is thus fundamentally different from demographic parity (Dwork et al., 2012; Yang and Stoyanovich, 2016) and affirmative action rules like Rooney rule (Collins, 2007; Celis et al., 2020), $\frac{4}{5}$th rule (selection rate for a protected group must be at least 80% of the rate for the group with the highest rate) [^3] Uniform Guidelines on Employment Selection Procedures, 29 C.F.R.§1607.4(D) (2015) or $\gamma$-based notions of fairness (Emelianov et al., 2020) and threshold based formulations such as FAIR (Zehlike et al., 2017).
@@
One of our key contributions includes formalizing the connection between $\delta$-EOR Fair Ranking described in Definition 4.1 and the cost of opportunity in rankings described in Section 4.1. Both $\delta$-EOR Fair Ranking and cost of opportunity in rankings are derived separately – the former from the axiom of fairness of a uniform lottery, the latter from the cost of errors that any realistic prediction model is bound to make. In the next section, we show that these two are elegantly related via theoretical results on cost optimality.
@@ 5. Computing EOR-Fair Rankings
Initialize: $j\leftarrow 0$; empty ranking $\sigma^{EOR}$ while $j<k$ do
       $l_{g}\leftarrow\sigma^{PRP,g}[1]\hskip 10.0pt\forall g\in\{A,B\}$
       $g^{*}\leftarrow\underset{g\in\{A,B\}}{\arg\min}\left|\delta(\sigma^{EOR}\cup\{l_{g}\})\right|$,
       where $\delta(.)$ is computed using (7)
       $l_{g^{*}}\leftarrow\sigma^{PRP,g^{*}}[1];\quad\sigma^{PRP,g^{*}}\leftarrow\sigma^{PRP,g^{*}}$$\{l_{g^{*}}\}$
       $\sigma^{EOR}\leftarrow\sigma^{EOR}\cup\{l_{g^{*}}\};\quad j\leftarrow j+1$
Return $\sigma^{EOR}$
@@
While Algorithm 1 is inspired by existing algorithms such as (Zehlike et al., 2017) in that both select the top element from the PRP ranking of each group, they are fundamentally different. Existing methods including (Zehlike et al., 2017) ensure a form of demographic parity which we have already shown to be fundamentally different than the EOR criterion we propose. Additionally, while (Zehlike et al., 2017) requires a threshold input and the designation of a disadvantaged group, the EOR Algorithm does not require this normative designation and guarantees EOR fairness without requiring any tolerance $\delta$ as an input. We show this both theoretically and in empirical evaluations and provide a detailed description of baseline algorithms in Appendix E.1.
@@
Minimize total cost as defined in Eq. (4)
(ILP) $\min_{x\in\{0,1\}}$ $\quad 1-\frac{P^{T}X}{nRel(A)+nRel(B)}$
(select up to $k$ candidates) s.t. $X^{T}\mathbb{1}=k\quad$
(EOR fairness from Eq. (6) must be satisfied $\forall k$) $-\delta(\sigma^{EOR}_{k})\leq$ $\left(\frac{P\mathbb{I}_{A}}{nRel(A)}-\frac{P\mathbb{I}_{B}}{nRel(B)}\right)^{T}X\leq\delta(\sigma^{EOR}_{k})$
We relax this ILP to a Linear Program (LP) by turning any integer constraints $x\in\{0,1\}$ in the primal into $0\leq x\leq 1$. For the relaxed LP, we formulate its dual and construct a set of dual variables $\lambda$ corresponding to the solution from the EOR Algorithm. Using the dual value of the EOR solution and the relaxed LP solution, we obtain an upper bound of the duality gap. Since the upper bound on this duality gap is w.r.t. the relaxed LP solution, it is also an upper bound for the optimal ILP solution. We provide a complete proof of the theorem and associated lemmas in Section C.1. ∎
@@
$\phi\delta_{max}=\frac{1}{nRel(A)+nRel(B)}\left|\frac{p_{A}-p_{B}}{q_{A}+q_{B}}\right|\left(\frac{\sigma^{PRP,A}[1]}{nRel(A)}+\frac{\sigma^{PRP,B}[1]}{nRel(B)}\right).$ Further, we show that if the EOR algorithm selects all the elements from one group at some position $k$, then selecting the remaining elements from the other group satisfies the $\delta_{max}$ constraint. We provide a complete proof of this theorem in Section C.2. ∎
We now compare EOR with the Uniform ranking policy and analyze positions $k$ with $\delta=0$ to avoid discretization effects.
@@
We provide the proof of Proposition 5.3 in Section C.3. In summary, we have shown that Algorithm 1 is an efficient algorithm that computes rankings close to the optimal solution, making it a promising candidate for practical use.